OpenStudy (anonymous):

Assume that the population of heights of male college students is approximately normally distributed with mean m of 70.63 inches and standard deviation s of 6.47 inches. A random sample of 88 heights is obtained. Show all work. (A) Find the mean and standard error of the distribution

OpenStudy (kropot72):

Standard error is a measure of a reasonable difference between a sample mean and the parent mean. In this case the standard error for a random sample of 88 heights is: \[\frac{\sigma}{\sqrt{n}}=\frac{6.47}{\sqrt{88}}=?inches\]

OpenStudy (anonymous):

already finsh this problem thanks